# Equivariant concordance of periodic 2-knots in $S^{4}$

### v1

@@ 4.1. Spin structures and ambient surgery in dimension 5
We take a moment to describe a restraint on ambient surgeries for 3-manifolds in dimension 5. We adapt the language of Sunukjian [Sun15], although this obstruction was known to Kervaire and overcome in his original proof of sliceness of 2-knots. Note that for this discussion, we do not assume any equivariance.
Recall that a trivialization $\xi$ of a vector bundle $E$ over base $M^{n}$ with fiber $F^{k}$ is an identification of $E$ with $M\times F$. One can specify a trivialization directly with a map, or with a choice of $k$ linearly independent sections of $E$. We use the terms trivialization and framing interchangeably.
@@
Recall that a spin structure on an $n$-manifold $M$ can be thought of as a trivialization of the tangent bundle $TM$ restricted to the 1-skeleton $M_{1}$ that extends over the 2-skeleton $M_{2}$. A discussion of the equivalence of this definition to the usual one involving lifts of principal $\text{SO}(n)$-bundles can be found in [Kir89, Ch. VI].
@@
Finally, we recall the definition of ambient surgery. Let $Y$ be a 3-manifold embedded in $B^{5}$, and let $\gamma$ be a curve in $Y$ such that $\gamma$ bounds a disk $\Delta$ into $B^{5}\setminus Y$. Let $E$ be the total space of an embedded 2-disk sub-bundle of the normal bundle $\nu_{B^{5}}(\Delta)$ such that $\partial E\cap Y=\nu_{Y}(\gamma)$. Finally, let $\xi$ be a framing of $\nu_{Y}(\gamma)$ that extends to the unique trivialization of $E$ over $\Delta$. Then the manifold $Y^{\prime}$ embedded in $B^{5}$ given by $Y^{\prime}:=Y\setminus\nu_{Y}(\gamma)\cup(\Delta\times\partial D^{2})$ is the result of ambient surgery along $\gamma$ with framing $\xi$.
@@
Let $Y$ be a 3-manifold embedded in $B^{5}$, and let $\gamma$ be a curve in $Y$ with a choice of framing $\xi$ such that $\gamma$ bounds a disk $\Delta$ into $B^{5}\setminus Y$. Let $\mathfrak{s}$ be the unique spin structure on $B^{5}$. Let $Y^{\prime}$ be the result of $\xi$-framed surgery along $\gamma$. Then the surgery $Y$ to $Y^{\prime}$ can be realized as an ambient surgery in $B^{5}$ if and only if $\xi$-framed surgery along $\gamma$ is a spin surgery with respect to $\mathfrak{s}|_{Y}$.
@@
Since $\gamma$ is a curve in an orientable a 3-manifold, trivializations of its tubular neighborhood $\nu_{Y}(\gamma)$ are in bijection with $\pi_{1}\text{SO}(2)\cong\mathbb{Z}$ and carry an action of $\mathbb{Z}$ which increases or decreases the number of full counterclockwise twists. On the other hand, trivializations of $\nu_{B^{5}}(\gamma)$ are in bijection with $\pi_{1}\text{SO}(4)\cong\mathbb{Z}/2\mathbb{Z}$. Indeed, the inclusion map of $\nu_{Y}(\gamma)$ into $\nu_{B^{5}}(\gamma)$ induces the mod 2 map from $\mathbb{Z}$ to $\mathbb{Z}/2\mathbb{Z}$. Thus, if $\xi$ and $\xi^{\prime}$ are two different trivializations of $\nu_{Y}(\gamma)$ that differ by an even number of twists, then when included into $\nu_{B^{5}}(\gamma)$, they become homotopic.
@@
Now $\Delta$ is embedded in $B^{5}$, and so the unique trivialization of its normal bundle $\nu_{B^{5}}(D)$ is exactly the unique spin structure $\mathfrak{s}$ on $B^{5}$ restricted to $\Delta$. This then further restricts to a trivialization $\zeta$ on any 2-disk sub-bundle of $\nu_{B^{5}}(\Delta)$, and in particular on $\nu_{B^{5}}(\partial\Delta)$, which is just $\nu_{B^{5}}(\gamma)$. Thus the framings $\xi$ on $\nu_{Y}(\gamma)$ for which there exist embedded $D^{2}$ sub-bundles of $\nu_{B^{5}}(\Delta)$ over which they can be extended are precisely those that agree with $\zeta$ restricted to $\nu_{B^{5}}(\gamma)$. Since $\zeta$ was induced by the restriction of the ambient spin structure on $B^{5}$, we see that surgery along $\gamma$ with framing $\xi$ must be spin surgery. ∎
4.2. Equivariant surgery preserving the fixed set
Here we address another restraint on the allowable surgeries that does not occur in non-equivariant surgery. Suppose we wished to surger along a curve $\gamma$ in $\overline{Y}$ that did not bound any disk disjoint from the image of $\operatorname{Branch}(\tau)$ in the quotient $\overline{B^{5}}$. This will certainly be the case if $\gamma$ lifts to a single curve, but even if it lifts to two curves, if these curves are not null-homotopic in $B^{5}\setminus(Y\cup\operatorname{Fix}(\tau))$, then when we try to surger $Y$ along them, we will introduce new connected components of the fixed set of our surgered 3-manifold. This is illustrated in Figure 8.
@@ 4.2. Equivariant surgery preserving the fixed set
Figure 8. New components of the fixed set introduced by errant handle attachment. On the left, we see $\gamma$ (horizontal) linking $\operatorname{Fix}(\tau)$ (vertical) inside $Y$. On the right, we see a different 3-dimensional projection of the same intersection, in which $\gamma$ appears as a pair of points and $\operatorname{Fix}(\tau)$ is a plane. The arc connecting these two points and intersecting $\operatorname{Fix}(\tau)$ is one slice of the disk $\Delta$ that we want to surger along, with its normal neighborhood shown in blue. However, as we see, this neighborhood intersects $\operatorname{Fix}(\tau)$, and so surgery along $\gamma$ will introduce a new component of the fixed set of $Y$ given by the intersection of the boundary of the tubular neighborhood of $\Delta$ with $\operatorname{Fix}(\tau)$.
It is a theorem of Birman-Hilden [BH75] together with Morgan’s work on the Smith Conjecture [Mor84] that if the cyclic branched cover of $S^{3}$ over any link is itself $S^{3}$, then the link is the unknot. Thus our only hope for surgering $Y$ to $B^{3}$ is to make sure our surgeries on $\overline{Y}$ do not introduce new components of $\overline{A}$. Therefore, in order to check if a curve $\gamma$ in $\overline{Y}$ can be surgered along, we need a way to check that it bounds a disk that does not intersect either $\overline{Y}$ or $\operatorname{Branch}(\tau)$. The following proposition will help us do this.

### v2

@@ 5.2.1. Spin structures and ambient surgery in dimension 5
We first describe a restraint on ambient surgeries for 3-manifolds in dimension 5. We adapt the language of Sunukjian [SUN15], although this obstruction was known to Kervaire and overcome in his original proof of sliceness of 2-knots.
@@
Recall that a trivialization $\xi$ of a vector bundle $E$ over base $M^{n}$ with fiber $F^{k}$ is an identification of $E$ with $M\times F$. One can specify a trivialization directly with a map, or with a choice of $k$ linearly independent sections of $E$. We use the terms trivialization and framing interchangeably. Recall also that a spin structure on an $n$-manifold $M$ can be thought of as a trivialization of the tangent bundle $TM$ restricted to the 1-skeleton $M_{1}$ that extends over the 2-skeleton $M_{2}$. A discussion of the equivalence of this definition to the usual one involving lifts of principal $\operatorname{SO}(n)$-bundles can be found in [KIR89, Ch. VI].
@@
Finally, we recall the definition of ambient surgery. Let $Y$ be a 3-manifold embedded in $B^{5}$, and let $\gamma$ be a curve in $Y$ such that $\gamma$ bounds a disk $\Delta$ into $B^{5}\setminus Y$. Let $E$ be the total space of an embedded 2-disk sub-bundle of the normal bundle $\nu_{B^{5}}(\Delta)$ such that $\partial E\cap Y=\nu_{Y}(\gamma)$. Lastly, let $\xi$ be a framing of $\nu_{Y}(\gamma)$ that extends to the unique trivialization of $E$ over $\Delta$. Then the manifold $Y^{\prime}$ embedded in $B^{5}$ given by $Y^{\prime}:=Y\setminus\nu_{Y}(\gamma)\cup(\Delta\times\partial D^{2})$ is the result of ambient surgery along $\gamma$ with framing $\xi$.
@@
Let $Y$ be a 3-manifold embedded in $B^{5}$, and let $\gamma$ be a curve in $Y$ with a choice of framing $\xi$ such that $\gamma$ bounds a disk $\Delta$ into $B^{5}\setminus Y$. Let $\mathfrak{s}_{0}$ be the unique spin structure on $B^{5}$. Let $Y^{\prime}$ be the result of $\xi$-framed surgery along $\gamma$. Then the surgery $Y$ to $Y^{\prime}$ can be realized as an ambient surgery in $B^{5}$ if and only if $\xi$-framed surgery along $\gamma$ is a spin surgery with respect to $\mathfrak{s}_{0}|_{Y}$.
@@
Since $\gamma$ is a curve in an orientable a 3-manifold, trivializations of its tubular neighborhood $\nu_{Y}(\gamma)$ are in bijection with $\pi_{1}\operatorname{SO}(2)\cong\mathbb{Z}$ and carry an action of $\mathbb{Z}$ which increases or decreases the number of full counterclockwise twists. On the other hand, trivializations of $\nu_{B^{5}}(\gamma)$ are in bijection with $\pi_{1}\operatorname{SO}(4)\cong\mathbb{Z}/2\mathbb{Z}$. Indeed, the inclusion map of $\nu_{Y}(\gamma)$ into $\nu_{B^{5}}(\gamma)$ induces the mod 2 map from $\mathbb{Z}$ to $\mathbb{Z}/2\mathbb{Z}$. Thus, if $\xi$ and $\xi^{\prime}$ are two different trivializations of $\nu_{Y}(\gamma)$ that differ by an even number of twists, then when included into $\nu_{B^{5}}(\gamma)$, they become homotopic.
@@
Now $\Delta$ is embedded in $B^{5}$, and so the unique trivialization of its normal bundle $\nu_{B^{5}}(D)$ is exactly the unique spin structure $\mathfrak{s_{0}}$ on $B^{5}$ restricted to $\Delta$. This then further restricts to a trivialization $\zeta$ on any 2-disk sub-bundle of $\nu_{B^{5}}(\Delta)$, and in particular on $\nu_{B^{5}}(\partial\Delta)$, which is just $\nu_{B^{5}}(\gamma)$. Thus the framings $\xi$ on $\nu_{Y}(\gamma)$ for which there exist embedded $D^{2}$ sub-bundles of $\nu_{B^{5}}(\Delta)$ over which they can be extended are precisely those that agree with $\zeta$ restricted to $\nu_{B^{5}}(\gamma)$. Since $\zeta$ was induced by the restriction of the ambient spin structure on $B^{5}$, we see that surgery along $\gamma$ with framing $\xi$ must be spin surgery. ∎
5.2.2. Avoiding the removed 3-ball
Here we address the restraint on the allowable surgeries imposed by requiring that a slice ball we construct for $S$ actually give a slice cylinder for the corresponding annulus $N$.
Suppose we wished to surger along a curve $\gamma$ in $Y$ bounding a disk $\Delta$ in $B^{5}$ such that $\Delta\cap Q$ was nonempty. Generically, $\Delta$ intersects $Q$ in points, so let’s examine one such intersection point $q$. When $Y$ is replaced by $Y^{\prime}=(Y\setminus\nu(\gamma))\cup_{\partial(\nu(\gamma))}\partial(H)$ for $H$ some 2-disk sub-bundle of $\nu(\Delta)$ thought of as a 4-dimensional 2-handle, $Y^{\prime}\cap Q$ now includes the belt sphere of $H$ over $q$. This is illustrated in Figures 13 and 14.
[…]
Figure 13. A schematic of two disks bounded by curves in $Y$. Note that the disk on the right is disjoint from $Q$, but the disk on the left intersects it transversely in a single point.
[…]
@@ 5.2.2. Avoiding the removed 3-ball
Figure 14. New components of $Y\cap Q$ introduced by errant handle attachment. On the left, we see an example surgery curve $\gamma$ linking $Q$ inside $Y$ (and thus linking their intersection, which is AA). On the right, we see a different 3-dimensional projection of the same intersection, in which $\gamma$ appears as a pair of points and $Q$ as a plane. The arc connecting these two points and intersecting $Q$ is one slice of the disk $\Delta$ that we want to surger along, with the 2-disk bundle $H$ shaded. However, as we see, this neighborhood intersects $Q$ in a disk, and so surgery along $\gamma$ will introduce a new component of $Y\cap Q$ given by the intersection of the boundary of $H$ with $Q$.
This is a major problem for our surgery program, because once a new closed component of this intersection is introduced, it is not clear that it can be removed by ambient surgery. We therefore want to restrict ourselves to surgery along curves in $Y$ which bound disks with interiors in $B^{5}\setminus(Y\cup Q)$. Thankfully, we can compute the fundamental group of this complement.
Proposition 5.8.
$\pi_{1}(B^{5}\setminus(Y\cup Q))\cong\mathbb{Z}^{2}$.
Proof.
We make a rising waters argument; compare with [GS00], Prop. 6.2.1 for this type of argument for surfaces in 4-manifolds. Consider the radial Morse function $f$ on $B^{5}$ from Lemma 5.3. Recall that both $Y$ and $Q$ are properly embedded and boundary parallel, and their intersection is a single properly embedded arc $A$. The key observation is that by construction, $Y,Q,$ and $A$ each have a single 0-handle in the handle decomposition induced by the restriction of $f$.
The 0-handles of $Y$ and $Q$ each contribute a 1-handle each to the complement. The minimum point of their intersection then contributes a 2-handle, the attaching sphere of which winds around the commutator of the meridians of both $Y$ and $Q$ near the intersection point. Thus the fundamental group of the sublevel set of $f$ just above the first intersection point is $\mathbb{Z}^{2}$. However, all subsequent critical points of $Y\cup Q$ are in $Y$ and are of index 1 or 2, which contribute 2- or 3-handles to the complement. Of these, only the 2-handles can contribute to the fundamental group relations of $B^{5}\setminus(Y\cup Q)$, but these only introduce extra commutation relations between meridians that already commute. ∎
Hence we conclude that if $\gamma$ is a curve in $Y$ which is null-homologous in $B^{5}\setminus Q$, then it bounds a disk in $B^{5}\setminus(Y\cup Q)$ and we may therefore surger along it (with a spin framing) without introducing new components of $Y\cap Q$. Note that this condition is satisfied it $\gamma$ is null-homologous in $Y\setminus A$.
5.2.3. Finding surgery curves
The discussions of the previous two subsubsections amount to the following rule when performing ambient surgery in this setup: we may surger $Y$ along any curve $\gamma$ with framing $\xi$ so long as:
  1. (1)
    $\xi$-framed surgery is spin surgery with respect to the restriction of the ambient spin structure $\mathfrak{s}_{0}$ on $B^{5}$ to $Y$, and
  2. (2)
    $\gamma$ must be null-homologous in $Y\setminus A$.
For ease of exposition, we will call pairs $(\gamma,\xi)$ satisfying these properties viable, and surgeries of $Y$ along them viable surgeries.
Our invariant $\alpha$ was chosen precisely so as to be invariant under viable surgeries:
Proposition 5.9.
Let $Y$ be a Seifert manifold for $S$ intersecting $Q$ in a null-homologous arc $A$, and let $Y^{\prime}$ be a 3-manifold properly embedded in $B^{5}$ bounded by $S$ intersecting $Q$ in null-homologous arc $A^{\prime}$. Suppose that $Y^{\prime}$ is related to $Y$ by viable surgeries. Then $\operatorname{Arf}(A,Y,\mathfrak{s}_{0})=\operatorname{Arf}(A^{\prime},Y^{\prime},\mathfrak{s}_{0})$.
Proof.
Recall that abstractly, we may cap off each pair of arcs and 3-manifolds with trivial arcs in a 3-ball, and that we use these closed versions to compute the Arf invariant, as in Definition 2.13. For ease of argument, continue to call the capped-off manifolds $Y$ and $Y^{\prime}$, and the knots inside them $A$ and $A^{\prime}$.
A collection of viable surgeries $(\gamma_{i},\xi_{i})$ relating $Y$ and $Y^{\prime}$ gives a recipe for a smooth characteristic concordance in a spin 4-manifold $X$ (in the sense of Definition 2.11) between $(A,Y,\mathfrak{s}_{0})$ and $(A^{\prime},Y^{\prime},\mathfrak{s}_{0})$. Specifically, because viable surgery curves are required to be null-homologous in $Y\setminus A$, the product $A\times I$ embeds in $X$ as a characteristic annulus cobounded by $A$ and $A^{\prime}$. Thus by Proposition 2.12, $\operatorname{Arf}(A,Y,\mathfrak{s}_{0})=\operatorname{Arf}(A^{\prime},Y^{\prime},\mathfrak{s}_{0})$. ∎
With these restrictions in mind, we turn to finding viable surgery curves and disks, which we do in two steps. The first is to ignore $A$ and simply find surgery curves that realize an embedded spin cobordism between $Y$ and $B^{3}$ inside $B^{5}$. The second is to then modify the 3-ball we surgered $Y$ to in order to unknot its intersection with $Q$. This is really the same procedure as the construction of spin characteristic fillings given in Proposition 2.9, but performed ambiently inside of $B^{5}$.
Lemma 5.10.
There exists a sequence of viable surgeries from $Y$ to $B^{3}$ such that the resulting $B^{3}$ intersects $Q$ in a single arc.
Proof.
Recall that by Kaplan’s theorem [KAP79], for any spin 3-manifold $(Y,\mathfrak{s})$, there exists a sequence of spin surgeries taking $(Y,\mathfrak{s})$ to the 3-ball with its unique spin structure $(B,\mathfrak{s}_{B})$. We use this to find starting curves and framings for our surgeries on $(Y,\mathfrak{s}_{0})$, since by Lemma 5.7 we can realize any spin surgery as an ambient surgery in $B^{5}$. Once we identify the curves we wish to surger $Y$ along, we isotope them inside $Y$ so that they are null-homologous in $B^{5}\setminus(Y\cup Q)$, allowing them to pass through $A$ if necessary. Such an isotopy inside $Y$ can always be found as $A$ is null-homologous in $Y$ rel. boundary, and so if a surgery curve $\gamma$ in $Y$ does not bound a disk in $B^{5}\setminus Y\cup Q$, we may use a relative Seifert surface for $A$ in $Y$ bounded by $A$ to isotope it along a path to $A$ and push it past $A$ to remove the intersection. This is similar to the trick shown in Figure 6. The result of the surgery on these curves will be to turn $Y$ into a 3-ball $B$, realizing $A$ as a knotted arc in $B$. ∎
It remains now to unknot $A$ through viable ambient surgeries. These surgeries must preserve the homeomorphism type of the 3-ball in which $A$ lives, and thus amount to changing the isotopy class of $A$ as an arc in $B^{3}$ by changing which $B^{3}$ in $B^{5}$ we think of $A$ as living in.
Lemma 5.11.
Suppose that $\operatorname{Arf}(A,B,\mathfrak{s}_{0})=0$. Then there exists a sequence of viable surgeries on $B$ which realizes $A$ as an unknotted arc in a 3-ball.
Proof.
Note that since $B$ is a 3-ball and therefore has a unique spin structure, $\operatorname{Arf}(A,B,\mathfrak{s}_{0})$ is equal to the classical Arf invariant of the knot $K$ given by capping off $B$ with another copy of $B^{3}$ and $A$ with a trivial arc. Recall by Lemma 2.10 that if $\operatorname{Arf}(K)=0$, then $K$ is band pass equivalent to the unknot, and that these band passes can be achieved by 0-framed (thus spin) surgery on curves which are null-homologous in the complement of $K$ in $S^{3}$, as shown in Figure 4. These curves and coefficients then represent viable surgeries on $B$. We may take these curves to be contained in $B$, and since they are null-homologous in $B\setminus A$, by Proposition 5.8, they bound disks in the complement of $B\cup Q$ in $B^{5}$, and hence surgering along them does not introduce any new components of $B\cap Q$. The result of these surgeries is then a (possibly different) 3-ball bounded by $S$ which intersects $Q$ in an unknotted arc $\ell$. ∎
We can now put together the entire ambient surgery argument to construct slice cylinders.
Proposition 5.12.
If $\alpha(N)=0$, then $N$ is slice.
Proof.
Let $N$ be a Montesinos annulus, let $S$ be its corresponding sphere, and let $R$ be its removed sphere. By Proposition 4.1 and Lemma 5.3, $S$ bounds a Seifert 3-manifold $Y$ which can be pushed into $B^{5}$ such that its intersection with $Q$ is an arc $A$. By Lemma 5.10, $Y$ can then be ambiently surgered to a copy $B$ of $B^{3}$ through a sequence of viable surgeries, thus preserving $Y\cap Q$.
By the definition of the invariant and Proposition 5.9, the assumption that $\alpha(N)=0$ implies that $\operatorname{Arf}(A,Y,\mathfrak{s}_{0})=0$. Therefore by Lemma 5.11, there exists a sequence of viable surgeries on $B$ which turn $B\cap Q$ into an unknotted arc $\ell$ in a 3-ball $B^{\prime}$. Since $B^{\prime}$ is a 3-ball bounded by $S$ with $B^{\prime}\cap Q=\ell$ an unknotted arc, by removing a neighborhood of $Q$, we obtain a slice cylinder for $N$. ∎
We may now show that the $\alpha$ invariant completely classifies Montesinos annuli up to concordance.
Proof of Theorem 5.2.
Together, Propositions 5.12 and 5.5 show that $N$ is slice if and only if $N$ has vanishing $\alpha$ invariant. We can then deduce that all annuli with nonvanishing $\alpha$ invariant are concordant by exploiting the group structure of annuli under stacking. If $N$ and $N^{\prime}$ are two Montesinos annuli such that $\alpha(N)=\alpha(N^{\prime})=1$, then $\alpha(N\natural N^{\prime})=0$ by Lemma 5.6, and so by the above construction, $N\natural N^{\prime}$ is slice. We therefore have that every annulus with nonvanishing $\alpha$ invariant is the inverse of every other annulus with nonvanishing $\alpha$ invariant, and so since these annuli cannot represent the identity element in $\mathcal{C}_{2}^{\circ}$ by Proposition 5.5, they must all belong to the same concordance class which has order 2 in the concordance group.
To see that $N_{T}$ generates $\mathcal{C}_{2}^{\circ}$, recall that in Example 1.5 we constructed a 3-ball bounded by $S^{T}$ which intersects the removed sphere in a trefoil tangle. Since the trefoil has Arf invariant 1, we have that $\alpha(N_{T})=1$, and so it cannot be slice by Proposition 5.5. ∎
6. Concordance of periodic 2-knots
To port over the classification given in Theorem 5.2 to the setting of periodic 2-knots, we take a moment to recall the relationship between periodic 2-knots and Montesinos annuli. Let $\tilde{S}$ be a $d$-periodic 2-knot in $S^{4}$. Then we may take the quotient of $\tilde{S}$ by $\rho$ to obtain a new sphere $S$ inside $S^{4}$ whose intersection with $\operatorname{Branch}(\rho)$ consists of exactly two points. Since $\operatorname{Branch}(\rho)$ is an unknotted 2-sphere, we may remove a neighborhood of it to obtain a Montesinos annulus $N$ in $S^{1}\times B^{3}$ given by $S\setminus\nu(\{p_{+},p_{-}\})$. We call the annulus $N$ obtained in this way the quotient annulus of $\tilde{S}$. The result we would like to prove is:
Theorem 6.1.
For each positive integer $d$, there exists a $\mathbb{Z}/2\mathbb{Z}$-valued, additive concordance invariant $\alpha$ for $d$-periodic 2-knots $\tilde{S}$ such that:
  1. (1)
    if $\alpha_{d}(\tilde{S})=0$, then $\tilde{S}$ is $\mathbb{Z}/d\mathbb{Z}$-equivariantly slice,
  2. (2)
    if $\alpha_{d}(\tilde{S})=1$, then $\tilde{S}$ is $\mathbb{Z}/d\mathbb{Z}$-equivariantly concordant to the $d$-twist spun trefoil, and
  3. (3)
    every $d$-periodic 2-knot satisfies exactly one of the above two conditions.
Thus the smooth equivariant concordance group $\mathcal{C}_{2}^{d}$ is isomorphic to $\mathbb{Z}/2\mathbb{Z}$ for all $d$.
To prove this, we will need to successively reprove many results from earlier in this paper in the equivariant setting by passing from equivariant objects to their quotients, applying earlier results, and then passing back by taking an appropriate cyclic branched cover.
6.1. Equivariant Seifert manifolds
The first step of this process is to construct equivariant Seifert manifolds for simply periodic 2-knots and 3-manifolds. The relationship between these results and those in Section 4 is simply that the ambient subspaces $R$ and $Q$ mentioned in those proofs here appear as the images of the fixed sets of $\rho$ under the quotient map. Recall our convention that objects which carry a cyclic group action are denoted with tildes, and objects which are invariant under the action are denoted with hats. Objects in the quotient have no decoration.
Proposition 6.2.
Let $\tilde{S}$ be a periodic 2-knot in $\tilde{S}^{4}$. Then $\tilde{S}$ bounds a $\rho$-equivariant Seifert 3-manifold $\tilde{Y}\subset\tilde{S}^{4}$ such that $\operatorname{Fix}(\rho)\cap\tilde{Y}$ consists of an arc $A$ that is null-homologous in $\tilde{Y}$ rel. boundary.
Proof.
Let $S$ be the quotient sphere for $\tilde{S}$ inside the copy of $S^{4}$ obtained after taking the quotient by $\rho$. Then $\operatorname{Branch}(\rho)$ is an unknotted 2-sphere intersecting $S$ in exactly two points. We may therefore apply Proposition 4.1 to obtain a Seifert 3-manifold $Y$ for $S$ that intersects $\operatorname{Branch}(\rho)$ in an arc $A$ which is null-homologous in $Y$ rel. boundary. Since $A$ is contained in $\operatorname{Branch}(\rho)$, after taking the $d$-fold cyclic branched cover of $S^{4}$ along $\operatorname{Branch}(\rho)$, $Y$ lifts to an equivariant Seifert 3-manifold $\tilde{Y}$ whose intersection with $\operatorname{Fix}(\rho)$ is the lift of $A$. Finally, since $A$ was null-homologous in $Y$ rel. boundary, its lift is null-homologous in $\tilde{Y}$ rel. boundary. ∎
The construction of equivariant Seifert 4-manifolds below is much the same.
Proposition 6.3.
Let $\tilde{Y}$ be a closed, simply periodic 3-manifold smoothly embedded in $\tilde{S}^{5}$, and suppose that $\tilde{Y}\cap\operatorname{Fix}(\rho)$ is a null-homologous knot $\hat{K}$ in $\tilde{Y}$. Suppose that there exists an equivariant Morse function $\tilde{f}$ on $\tilde{S}^{5}$ such that $\tilde{Y}$ is in ordered equivariant Morse position with respect to $\tilde{f}$, $\tilde{f}|_{\hat{K}}$ has exactly two critical points, and away from these two critical points, $\hat{K}$ is fully vertical with respect to $\tilde{f}|_{\tilde{Y}}$. Then $\tilde{Y}$ bounds a $\rho$-equivariant Seifert 4-manifold $\tilde{X}$ smoothly embedded in $\tilde{S}^{5}$ such that $\tilde{X}\cap\operatorname{Fix}(\rho)$ is a smoothly embedded disk $\hat{D}$ bounded by $\hat{K}$ which is null-homologous in $\tilde{X}$ rel. boundary.
Proof.
Let $Y$ be the quotient of $\tilde{Y}$ under $\rho$ inside $S^{5}$ such that $Y\cap\operatorname{Branch}(\rho)$ is the knot $K$ given by the image of $\hat{K}$ under the covering map. Since $\tilde{f}$ is equivariant, it also descends to a Morse function $f\colon S^{5}\to\mathbb{R}$. Since $\tilde{Y}$ was in ordered equivariant Morse position with respect to $\tilde{f}$, $Y$ is in ordered Morse position with respect to $f$, and since $\hat{K}$ was contained in $\operatorname{Fix}(\rho)$, $f|_{K}$ has exactly two critical points and $K$ is otherwise fully vertical with respect to $f$. We therefore can apply Proposition 4.2 to obtain a Seifert 4-manifold $X$ for $Y$ smoothly embedded in $S^{5}$ such that its intersection with $\operatorname{Branch}(\rho)$ is a smoothly embedded disk $D$ bounded by $K$ which is null-homologous in $X$ rel. boundary. Taking the branched cover of the whole setup over $\operatorname{Branch}(\rho)$ then recovers $\tilde{Y}$ with $\tilde{Y}\cap\operatorname{Fix}(\rho)=\hat{K}$ and the lifts of $X$ and $D$ give $\tilde{X}$ and $\hat{D}$ with the desired properties. ∎
6.2. The equivariant concordance invariant
This subsection reproves the contents of Subsection 5.1 in the context of periodic 2-knots. For the following, fix a period $d$ for $\rho$.
Definition 6.4.
Let $\tilde{S}$ be a $d$-periodic 2-knot in $\tilde{S}^{4}$. Let $\tilde{Y}$ be an equivariant Seifert 3-manifold for $\tilde{S}$ with $\tilde{Y}\cap\operatorname{Fix}(\rho)=\hat{A}$ a null-homologous arc in $\tilde{Y}$. Consider the manifold $Y$ and contained arc $A$ given by taking the quotient by $\rho$. We define the invariant $\alpha_{d}$ of $\tilde{S}$ to be:
$\alpha_{d}(\tilde{S}):=\operatorname{Arf}(A,Y,\mathfrak{s}_{0})\in\mathbb{Z}/2\mathbb{Z},$
where $\mathfrak{s}_{0}$ denotes the spin structure on $Y$ induced by the unique spin structure on $B^{5}$ via inclusion.